OpenStudy (anonymous):

Someone please help I dont know nothing about this answers in increasing order and r educed fraction form 3 b 2 +22b− 16 = 0

OpenStudy (mathmate):

To find the solution set, factorize the equation as:\ (b+8)(3b-2)=0 which gives the solution set as S={-8, 2/3} increasing order means that you would put the solution set from smallest root to the greatest. {-8,2/3} would be in increasing order, while {2/3,-8} would be in decreasing order. reduced fraction form means that 2/3 is in reduced fraction form. Even though 4/6 is an equivalent fraction, it is not in reduced fraction form. Hope that clears it up.

OpenStudy (anonymous):

thanks so much im having a hard time tryin to pass my final and i really dont remember algebra at all and i only have 1hr left to finish so thanks for the help

OpenStudy (anonymous):

3x + 4 ------ 2 = 8 can u help with this one, in increasing order and r educed fraction form

OpenStudy (mathmate):

There is only one solution in the solution set, which is: (3x+4)=2*8 3x+4=16 3x=12 x=4 Since the solution is a single integer, it is both in increasing order and in reduced fraction form.

OpenStudy (anonymous):

can u do this for me....Write an equation of the line with the given properties. Your answer should be written in slope-intercept form, if possible. Passing through (9, −7); m = 5